OpenStudy (elyssac):

PLEASE HELP!!! Find all solutions: 3cos[(pi/5)x]=2 Please help

OpenStudy (anonymous):

I'm sorry I don't understand this :/ I will google it and assist you upon your search! :)

OpenStudy (elyssac):

The answers apparently are: 0.1339+10k and 8.6614+10k where k is an integer. The problem is, I don't know how to SOLVE to get the answers. They give us a select few answers, but I still need to know the process of how to get there, if that makes sense.

OpenStudy (anonymous):

Alright I think I know the answer to your problem... Maybe Try this https://www.khanacademy.org/math/differential-calculus/limits_topic/algebraic-limits/v/limit-example-1 If this doesn't help I'll try to find more links to help you out :)

OpenStudy (anonymous):

Try this one love I think it fits more along the lines of what we're doing here. :3 https://www.youtube.com/watch?v=tqXQKTTlRWE

OpenStudy (elyssac):

Okay, I'll look at it and see if it's what I need. Thanks!

OpenStudy (anonymous):

Welcome Sorry I'm not that great at explaining or helping I try!

OpenStudy (baru):

\(cos(\frac{\pi x}{5})=3/2\\cos^{-1}(3/2)= \frac{\pi x}{5} +2k\pi\)

OpenStudy (baru):

sorry,typo , all 3/2 should be 2/3